MuCao Creamery has recently expanded into 3 provinces in Southeast China. Currently, there are three other retail ice cream vendors in these provinces along with MuCao with the following sales volumes:

Brands Total
Revenues
(RMB)
Cones
Served
(units)
Xihuan Cones 87,500 35,000
MuCao Creamery 110,000 50,000
Scoops 30,000 10,300
Swirls 59,200 20,000

1.What is MuCa's average price per cone?

2.What is the total market size in terms of revenues?
Answer #1

1.

Muca's average price per cone = 110000/50000 = 2.2 RMB

2.

Total market size in terms of revenues = 87500+110000+30000+59200 = 286700 RMB